OpenStudy (teamrockettwin):

Let f(x) = x − 2 and g(x) = x2 − 7x − 9. Find f(g(−1)). −21 −3 3 21

OpenStudy (herpderp):

Let f(x) = x − 2 and g(x) = x2 − 7x − 9. Find f(g(−1)). g(-1)=-1^2-7(-1)-9 g(-1)=1 + 8 - 9 g(-1)= 0 f(g(-1)) = 0 - 2 = -2 Are those answer choices correct

OpenStudy (teamrockettwin):

Those are the only answer choices they gave me :(

OpenStudy (herpderp):

Aw, maybe I did something wrong then O_o

OpenStudy (herpderp):

Wait I did

OpenStudy (herpderp):

My bad

OpenStudy (herpderp):

g(-1)=-1^2-7(-1)-9 g(-1)=1 + 7 - 9 g(-1)= -1 f(g(-1)) = -1 - 2 = -3

OpenStudy (teamrockettwin):

Lol I chose -3 because it was the closest to your answer and I got a 100 on the test. So thanks for your help on both questions :)

OpenStudy (herpderp):

There it is xD Sorry